OpenStudy (anonymous):

The length of a rectangle is 5 centimeters less than twice its width. The perimeter of the rectangle is 26 cm. What are the dimensions of the rectangle? A. length = 5 cm; width = 5 cm B. length = 7 cm; width = 6 cm C. length = 6 cm; width = 7 cm D. length = 4 cm; width = 9 cm

OpenStudy (lilg132):

B 6 cm

OpenStudy (lilg132):

B. length = 7cm; width = 6cm

OpenStudy (lilg132):

do you want to know how to work it out?

OpenStudy (lilg132):

dont forget to click good answer :)

OpenStudy (anonymous):

The perimeter of a rectangle is the sum of the length of all the sides. By definition a rectangle has 4 sides, and has two pairs of sides of equal length (generally, call the long side the length and the short side the width). You can call the length "L" and the width "w". Therefore, the perimeter of any rectangle is P = L + W + L + W or: P = 2L+2W Now, in this question, the length is 5 less than the twice the width. This means the length = 2 * width - 5 or L = 2W - 5 It is given that the Perimeter (P) = 26 Substituting this information in the Perimeter equation gives: 26 = 2 ( 2W-5) + 2W Distribute the 2: 26 = 2*2W - 2*5 + 2W Combine the W's 26 = 6W - 10 Add 10 to both sides 36 = 6W Divide by 4 on both sides W = 6 Remember that L = 2W - 5, so substitute this value of W to get L: L = 2*6 - 5 L = 7 So the answer is B.

OpenStudy (anonymous):